An Earth-shattering Work Significantly Advancing the State of the Art

by

An Author, BCompSc

Thesis

Submitted by An Author

for fulfillment of the Requirements for the Degree of

Doctor of Philosophy (0190)

Supervisor: Dr. Joe Bloggs Associate Supervisor: Dr. Fred Nurke

Clayton School of Information Technology

Monash University

October, 2005 \begin{matrix}c \\ ◯ \end{matrix} Copyright

by An Author I luv youse all

iii

Contents

List of Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . v List of Figures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . vi Abstract . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . vii Acknowledgments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ix 1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Literature Review . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3 Figures and Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3.1 Figures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.2 Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3.2.1 Referencing test . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 4 Method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Appendix A This appendix should get a letter . . . . . . . . . . . . . . . . 9 Appendix B Simulation Source Code . . . . . . . . . . . . . . . . . . . . . . . 11 Vita . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 References . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Last Thing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

List of Tables

3.1 An example of a table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

List of Figures

3.1 An example of a figure. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

An Earth-shattering Work Significantly Advancing the State of the Art

An Author, BCompSc

                        authora@csse.monash.edu.au

Monash University, 2005

Supervisor: Dr. Joe Bloggs

                        bloggsj@csse.monash.edu.au

Associate Supervisor: Dr. Fred Nurke

                         nurkef@csse.monash.edu.au

Abstract

This thesis introduces … Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet.

Duis autem vel eum iriure dolor in hendrerit in vulputate velit esse molestie consequat, vel illum dolore eu feugiat nulla facilisis at vero eros et accumsan et iusto odio dignissim qui blandit praesent luptatum zzril delenit augue duis dolore te feugait nulla facilisi. Lorem ipsum dolor sit amet, consectetuer adipiscing elit, sed diam nonummy nibh euismod tincidunt ut laoreet dolore magna aliquam erat volutpat.

An Earth-shattering Work Significantly Advancing the State of the Art

Declaration

I declare that this thesis is my own work and has not been submitted in any form for another degree or diploma at any university or other institute of tertiary education. Information derived from the published and unpublished work of others has been acknowledged in the text and a list of references is given.

An Author

October 18, 2005

Acknowledgments

I would like to thank everyone who helped to make this possible. It has been an incredible journey of self-discovery, and I love every last one of you…

An Author

Monash University October 2005 x

Chapter 1

Introduction

Here I introduce the research area and my contributions to it.

Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet.

Duis autem vel eum iriure dolor in hendrerit in vulputate velit esse molestie consequat, vel illum dolore eu feugiat nulla facilisis at vero eros et accumsan et iusto odio dignissim qui blandit praesent luptatum zzril delenit augue duis dolore te feugait nulla facilisi. Lorem ipsum dolor sit amet, consectetuer adipiscing elit, sed diam nonummy nibh euismod tincidunt ut laoreet dolore magna aliquam erat volutpat.

Ut wisi enim ad minim veniam, quis nostrud exerci tation ullamcorper suscipit lobortis nisl ut aliquip ex ea commodo consequat. Duis autem vel eum iriure dolor in hendrerit in vulputate velit esse molestie consequat, vel illum dolore eu feugiat nulla facilisis at vero eros et accumsan et iusto odio dignissim qui blandit praesent luptatum zzril delenit augue duis dolore te feugait nulla facilisi.

Nam liber tempor cum soluta nobis eleifend option congue nihil imperdiet doming id quod mazim placerat facer possim assum. Lorem ipsum dolor sit amet, consectetuer adipiscing elit, sed diam nonummy nibh euismod tincidunt ut laoreet dolore magna aliquam erat volutpat. Ut wisi enim ad minim veniam, quis nostrud exerci tation ullamcorper suscipit lobortis nisl ut aliquip ex ea commodo consequat.

Duis autem vel eum iriure dolor in hendrerit in vulputate velit esse molestie consequat, vel illum dolore eu feugiat nulla facilisis.

At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, At accusam aliquyam diam diam dolore dolores duo eirmod eos erat, et nonumy sed tempor et et invidunt justo labore Stet clita ea et gubergren, kasd magna no rebum. sanctus sea sed takimata ut vero voluptua.

2 CHAPTER 1. INTRODUCTION est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat.

Literature Review

In this chapter I will demonstrate some of the extended citation capabilities provided by the Harvard package (Williams and Schnier, 1994). As well as supporting the standard L^{A}TEX \cite command, it provides a few other very useful ones.

The \cite command is best used when placing a citation at the end of sentence or phrase (as above). When you want to refer to the authors of a particular work, typically at the start of a sentence, \cite is not appropriate. This is particularly so if you are using a numerical or symbolic citation style. You should not start a sentence with

[2] says that this is most certainly …

In such situations you really need to give the authors’ names. The Harvard package provides a new command \citeasnoun, which allows you to produce things like:

Ade (1983) describes a means by which textures may be characterized … another approach is given in De Bonet and Viola (1998).

Atick, Griffin and Redlich (1996) note that humans have little or no difficulty in perceiving shape, yet find it extremely difficult to describe what they perceive.

Another useful new command provided by the Harvard package is \possessivecite. This is used when you want to use the authors in a possessive sense, as below:

Atick et al.’s (1996) experiments with hundreds of three-dimensional scans of human heads suggests that …

Note that an abbreviated version of the authors’ names has been used above. The Harvard package does this automatically after the first citation. This behaviour can be overridden if desired.

Note: The standard Harvard package is incompatible with pdflatex (via the html package). If you want to use the Harvard package with pdflatex, make sure you use the version supplied here (2.0.5a), which has been hacked to remove the problem.

In the long run, natbib is probably the better choice. Watch this space...

4 CHAPTER 2. LITERATURE REVIEW

Chapter 3

Figures and Tables

Here we will test that references to figures and tables work correctly.

3.1 Figures

Figure 3.1: An example of a figure.

See Figure 3.1.

3.2 Tables

See Table 3.1.

3.2.1 Referencing test

See Table 3.1 and Figure 3.1.

23121 1212 232 cat frog dog Table 3.1: An example of a table 6 CHAPTER 3. FIGURES AND TABLES

Method

Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet.

Duis autem vel eum iriure dolor in hendrerit in vulputate velit esse molestie consequat, vel illum dolore eu feugiat nulla facilisis at vero eros et accumsan et iusto odio dignissim qui blandit praesent luptatum zzril delenit augue duis dolore te feugait nulla facilisi. Lorem ipsum dolor sit amet, consectetuer adipiscing elit, sed diam nonummy nibh euismod tincidunt ut laoreet dolore magna aliquam erat volutpat.

Ut wisi enim ad minim veniam, quis nostrud exerci tation ullamcorper suscipit lobortis nisl ut aliquip ex ea commodo consequat. Duis autem vel eum iriure dolor in hendrerit in vulputate velit esse molestie consequat, vel illum dolore eu feugiat nulla facilisis at vero eros et accumsan et iusto odio dignissim qui blandit praesent luptatum zzril delenit augue duis dolore te feugait nulla facilisi.

Nam liber tempor cum soluta nobis eleifend option congue nihil imperdiet doming id quod mazim placerat facer possim assum. Lorem ipsum dolor sit amet, consectetuer adipiscing elit, sed diam nonummy nibh euismod tincidunt ut laoreet dolore magna aliquam erat volutpat. Ut wisi enim ad minim veniam, quis nostrud exerci tation ullamcorper suscipit lobortis nisl ut aliquip ex ea commodo consequat.

Duis autem vel eum iriure dolor in hendrerit in vulputate velit esse molestie consequat, vel illum dolore eu feugiat nulla facilisis.

At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, At accusam aliquyam diam diam dolore dolores duo eirmod eos erat, et nonumy sed tempor et et invidunt justo labore Stet clita ea et gubergren, kasd magna no rebum. sanctus sea sed takimata ut vero voluptua. est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat.

Consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet.

Appendix A

This appendix should get a letter

An appendix before the backmatter gets an automatically generated letter by which it can be referred to. This is Appendix A.

10 APPENDIX A. THIS APPENDIX SHOULD GET A LETTER

Appendix B

Simulation Source Code

You may want to investigate the lgrind program and package if you wish to include source code in your thesis 12 APPENDIX B. SIMULATION SOURCE CODE

Vita

Publications arising from this thesis include: Author, A. and Bloggs, J. (2002), A really catchy title. In The 31st International

Conference on Non-specific Computing. Capital City, Country. Bloggs, J. and Author , A. (2002), A very much longer and significantly less catchy

title. in Workshop on A Research Area. Springfield, USA.

Permanent Address: Clayton School of Information Technology Monash University Australia

This thesis was typeset with L^{A}TEX 2ε^{1} by the author.

1L^{A}TEX 2ε is an extension of L^{A}TEX. L^{A}TEX is a collection of macros for TEX. TEX is a trademark of the American Mathematical Society. The macros used in formatting this thesis were written by Glenn Maughan and modified by Dean Thompson and David Squire of Monash University.

14 VITA

References

Ade, F. (1983). Characterisation of textures by “eigenfilters”, Signal Processing 5: 451– 457.

Atick, J. J., Griffin, P. A. and Redlich, A. N. (1996). The vocabulary of shape: principal shapes for probing perception and neural response, Neural Computation 7(1): 1–5.

De Bonet, J. S. and Viola, P. (1998). Texture recognition using a non-parametric multi-scale statistical model, Proceedings of the 1998 IEEE Conference on Computer Vision and Pattern Recognition (CVPR’98), Santa Barbara, California, USA.

http://www.ai.mit.edu/~jsd/research/publications/1998/DeBonet-CVPR98.pdf

Williams, P. and Schnier, T. (1994). The Harvard family of bibliography styles, L^{A}TEX2e package documentation. (last accessed February 7, 2004).

http://www.ctan.org/tex-archive/macros/latex/contrib/harvard/

16 REFERENCES

Last Thing

This sort of appendix has no letter.